### v2

@@ Abstract
Foundation Models, which leverage large neural networks pre-trained on unlabelled data before fine-tuning for specific tasks, are increasingly being applied to specialised domains. Recent examples include ClimaX for climate and Clay for satellite Earth observation, but a Foundation Model for Space Object Behavioural Analysis has not yet been developed. As orbital populations grow, automated methods for characterising space object behaviour are crucial for space safety. Here, we present a Space Safety and Sustainability Foundation Model focusing on space object behavioural analysis using light curves. To build our Foundation Model, we implemented a Perceiver-Variational Autoencoder (VAE) architecture, pre-trained with self-supervised reconstruction and masked reconstruction on 227,000 light curves from the MMT-9 observatory. The VAE enables anomaly detection, space object motion prediction, and generation of synthetic light curves. We fine-tuned the model for anomaly detection & motion prediction using two independent light curve simulators (CASSANDRA and GRIAL respectively), using CAD models of boxwing, Sentinel-3, SMOS, and Starlink platforms. Our pre-trained model achieved a reconstruction mean squared error of 0.009, identifying potentially anomalous light curves through reconstruction difficulty. After fine-tuning, the model scored 88% and 82% accuracy, with 0.90 and 0.95 ROC AUC scores respectively in both anomaly detection and motion mode prediction (e.g., sun-pointing, spin, tumbling etc.). Analysis of high-confidence anomaly predictions on real data revealed distinct patterns including characteristic object profiles and satellite glinting. The motion mode prediction model successfully differentiated between various movement behaviours such as sun-pointing, spin, and tumbling. Our work demonstrates how self-supervised learning can simultaneously enable anomaly detection, motion prediction, and synthetic data generation from rich representations learned in pre-training. More broadly, our work supports space safety and sustainability through automated monitoring and simulation capabilities.
Keywords Space Object Behavioural Analysis, Space Foundation Model, Self-Supervised Learning, Light curve anomaly detection, Space Situational Awareness (SSA), Space Foundation Models, Generative AI
@@ 1 Introduction
Figure 1: Graphical structure of the paper. First, in Section 2.1 we describe pre-training our Foundation Model with a large unlabelled dataset of real light curves. We encode these into a rich latent representation with a Perceiver-VAE architecture, updating these representations based on three self supervised learning (SSL) tasks: Reconstruction, Forecasting, and Masking. We next analyse the results of pre-training, which includes pre-flagging of anomalies based on reconstruction difficulty (Sections 2.2-2.3, and forecasting quality of the model (2.4. Following this, we descibe fine-tuning our rich representations for two downstream tasks: a) anomaly prediction (Section 2.5, and b) motion prediction (Section 2.6). Finally, we demonstrate further utility of our representations by generating de novo datasets according to a particular motion type (Section 2.7).
@@
However, there is little current exploration on the best neural network architectures and training strategies for FMs for more specialised datasets/applications, and to our knowledge this is yet to be done for Space Domain Awareness (SDA). Of principal interest in SDA is anomaly detection, i.e., detecting space objects that are behaving in an unusual way, such as motions, manoeuvres, or those that are performing atypical functions. Characterising these atypical behaviours can be done through simulation of space object observations, but this is challenging, due to imperfect physical models and computational cost. Therefore, a Foundation model for SDA should be able to perform anomaly detection, integrating real observations with physical models in a computationally efficient way.
@@
Whilst VAEs offer an effective anomaly detection mechanism, they still need integration into a broader neural network framework to effectively process complex, high dimensional space object data. Transformer architectures are an attractive choice here, due to their ability to capture long range dependencies in sequential data. When choosing a Transformer architecture, there are numerous variants with respective strengths and weaknesses. For example, conventional transformer-based neural networks perform the computationally costly full self attention mechanism between all the inputs. This means that standard transformers scale poorly with larger dataset sizes. To address this limitation, DeepMind recently developed the Perceiver/Perceiver-IO, a transformer-based neural network architecture. [8, 11]. For a SDA Foundation Model, Perceiver has two main motifs that make it an attractive choice of architecture for training a FM (Figure 2), computational efficiency and heterogenous learning. Perceiver only performs the full self-attention mechanism in the latent (hidden) space which is computationally efficient as the latent vector is constrained in size. Additionally, as the substantive part of the learning of Perceiver is done within the latent space, this readily enables heterogenous learning, as the modality of the data has already been abstracted into the lower dimensional latent space.
Given the potential of VAEs as anomaly detectors, and the computational efficiency of Perceiver, a VAE-based Perceiver model is a promising approach for a computationally efficient, SDA FM for anomaly detection. In this work, we will use visible light curves as a test-bed for development of an SDA Foundation Model, due to the availability of large amounts of real observational data, and the maturity of light curve simulators. Here, we use the largest publicly released dataset of light curves, from the Mini-Mega TORTORA (MMT-9) observatory in Russia [12].
@@ 2 Results & Discussion
Using a held-out test portion of the MMT dataset, we examined the highest and lowest reconstruction error samples (Tables 1-2). Among the highest reconstruction error test samples (Table 1), we found a mixture of satellites and rocket bodies, some appearing frequently in the pre-training data. For example, light curves from EGS (AJISAI) occur 1064 times in the MMT dataset, yet four light curves appear in the ten highest reconstruction error. Similarly, light curves from GLOBSTAR M092 and CZ-3 R/B (occurring 285 and 269 times respectively) also showed high reconstruction errors. This challenges the intuitive assumption that objects appearing more frequently would be easier for the model to reconstruct. One possible explanation is that these light curves represent atypical behaviour for those particular space objects. For the lowest reconstruction error samples (Table 2), we again find a mixture of satellites and rocket bodies, with one example of debris. Unlike the highest reconstruction errors, these typically had fewer appearances in the pre-training data (ranging from 9-153).
Table 1: The ten highest reconstruction errors from the pre-flagged light curve samples contained in the test set. We evaluated the lowest-loss light curve model from Section 2.2 against the entire test dataset unseen during training/validation, recording the reconstruction error for each sample. Examination of the ten lowest reconstruction error light curves shows a variety of space objects (SOs). Seven light curves are from satellites (all of which are in orbit), and three belong to rocket bodies. US: United States of America, IND: India, PRC: China.
@@ 2.5 Fine-tuning for anomaly detection
Obtaining real observational data of satellites with known anomalies is extremely challenging, as the full set of possible anomalies is not well understood or catalogued, and organisations are opaque about anomalous space object behaviour for national security/commercial reasons. To address this data gap, we produced a synthetic dataset for fine-tuning using the framework CASSANDRA (Computational Agent For Space Situational Awareness And Debris Remediation Automation). Within this framework is an orbital simulator first introduced by Vasile et al. (2023), generating the light curves using a CAD model of an orbiting Space Object with a specific geometry. This dataset contains 800 simulations of four different satellite platforms: a boxwing satellite (e.g., Jason-3), Sentinel-3, SMOS, and Starlink. These space objects (SOs) orbit an observing ground station, with both light and spectral curves recorded as time courses. Within these 800 simulations, a subset contain anomalous events of varying magnitudes - specifically debris collisions that result in detectable changes in signal. These changes in signal are shown in Figure 6 F, H, where otherwise linear light curves abruptly change to a different profile, oscillating in magnitude between very high and low minima. Our fine-tuning approach combines the pre-training objective with supervised learning to better characterise space object behaviour (see Equation 8 for details). Briefly, we impose the additional constraint of classification, where our Perceiver-VAE is simultaneously maintaining reconstruction whilst learning to distinguish between anomalous and non-anomalous curves.
@@
We focus on light curves due to public availability of the MMT-9 real observational dataset, however the CASSANDRA simulator also generates hyperspectral light curves. Therefore, we also present finetuning results on normal/anomalous hyperspectra, allowing us to explore the versatility of our approach and model performance across a different data modality.
To understand stability of fine-tuning, we fine-tuned our pre-trained Perceiver-VAE five times to convergence on the light curve and hyperspectral curves from the CASSANDRA fine-tuning dataset (Figure 7). We then average training runs and compute the variance. For light curves, we found our pre-trained model could classify anomalies in the fine-tuning validation dataset with approximately 88% accuracy, and 0.9 Receiver Operating Characteristic Area Under the Curve (ROC AUC) (Figure 7 A-D). Throughout training, the correlation between reconstruction loss and classification accuracy increased to approximately 0.28 (Figure 7 E). This suggests that the features learned during reconstruction are useful in classifying anomalous light curves from non-anomalous. For the hyperspectra, we found similarly high accuracy after fine-tuning (Figure 7), with validation accuracies of approximately 84% and ROC AUC scores of approximately 0.86.
Figure 7: Fine-tuning results for anomaly detection, averaged across five training runs comparing pre-trained light curve Perceiver-VAE (A-E), and the pre-trained hyperspectra Perceiver-VAE (F-J). Both training and validation classification accuracy and Receiver Operating Characteristic (ROC) Area Under the Curve (AUC) increase over time for both data modalities (A-D), (F-I). Validation accuracies of approximately 88% and 84% are achieved for the light curve and hyperspectral models respectively, with narrow variance between runs (B, G). Additionally, validation ROC-AUC values of approximately 0.9 and 0.88 are achieved for the light curve model and hyperspectral model respectively (D, I), suggesting good discriminatory power between anomalous/normal curves. We also tracked the correlation between reconstruction loss and classification loss, finding that for the light curve fine-tuning this increased over time to approximately 0.28 (E), whereas for hyperspectral fine-tuning the opposite was true, where the correlation decreased over time to approximately 0.05 (J).
However, unlike the light curve fine-tuning, the correlation between reconstruction loss and classification accuracy decreased from 0.3 to 0.05 (Figure 7). This suggests that while the model initially utilised the reconstruction representations to assist classification, it later learned task-specific features for classification, effectively separating out the pre-training task from that of the fine-tuning. The variance in this correlation was notably higher during hyperspectra fine-tuning when compared to the light curve fine-tuning.
@@
Having fine-tuned our model to high classification accuracy, we next analysed the held-out test dataset, classifying test samples as anomalous/non-anomalous and visualising the top twelve most confidently predicted anomalies (Figure 8). We found a variety of SOs in this set, including several satellites (Figure 8 A, B, E, H, J, L). All of these satellites except the light curve belonging to FLOCK 4P 6 (Figure 8 E) exhibited ’J’ shaped curves, which simulations suggest may represent glinting from a highly reflective surface aligning with the sun [20]. We also detected light curves belonging to debris (Figure 8 C, D, G), all displaying prominent ’U’ shaped profiles i.e., where the minima in magnitude is approximately in the middle of the observation. Finally, we identified three other distinct light curve motifs (Figure 8 red dotted outlines): abrupt changes in magnitude which presented as peaks (Figure 8 E), step changes (Figure 8 F), or highly periodic variation (Figure 8 K).
@@ 2.6 Fine-tuning for motion prediction
We also fine-tuned on a separate dataset with ground-truth motions, simulated by the GRIAL simulator from GMV [21]. Example light curves from this dataset are visualised in Figure 9, here we briefly summarise the dataset. The GMV dataset is a simulated dataset of 22,009 light curves of Sentinel-3A simulated under 10 different motion laws as seen from 30 different ground stations. These motions can be grouped into three distinct behaviours: First, Sun oriented: SafeX, SafeZ (where the Space Object’s X/Z axis points to the Sun respectively, and Y initially points to the celestial North pole, and the X/Z rotates respectively), and Sun (X-axis points to Sun, Y-axis to Celestial North pole, differentiated from Safe by a varied X-axis phase angle. Next, Earth oriented motions: YAWXC, YAWZC, YAWXS, YAWZS (where X/Z indicates the Space Object’s nadir pointing axis, and C/S indicates whether the object’s motion is compensating (C) for optical distortion, or (S) maximising solar array lighting). Finally, the dataset also contains labelled general motions less related to satellite function: Tumbling (uncontrolled/complex rotation of the SO), Spin (rotation around a single axis), Inertial pointing (where the SO has a fixed orientation relative to the J2000 reference frame). To fine-tune our pre-trained Perceiver-VAE, we first group similar motions together, as domain expert manual inspection between e.g., different yaw axial motions is a substantially more challenging task than distinguishing between different motions altogether (e.g., tumbling/sun pointing). After fine-tuning for 100 epochs, our motion classification tuned Foundation model could classify held out light curves with 82.74% accuracy, with an average ROC-AUC of 0.95 (Table 3). There was inter-class variation in the fine-tuned classifier’s performance. For example, both tumbling and spin show high performance, with precision of approximately 0.97 and 0.90 respectively. Similarly, these classes also showed high recall, both with approximately 0.93. In contrast, inertial has high precision (0.86) but low recall (0.33).
@@
Table 3: Performance evaluation of the motion classification model on a held-out test dataset. The model achieves 82.74% overall accuracy with an average ROC AUC of 0.9512. The metrics demonstrate varied performance across different motion classes, with TUMBLING and SPIN showing the strongest classification performance, while INERTIAL exhibits high precision but low recall. For YAW, the most common class, the model shows high recall but lower precision.
Motion Class Precision Recall F1-Score ROC AUC n
INERTIAL 0.8600 0.3282 0.4751 0.8686 262
SAFE 0.7711 0.7485 0.7596 0.9480 342
SPIN 0.9009 0.9289 0.9147 0.9912 225
SUN 0.6761 0.6316 0.6531 0.9454 152
TUMBLING 0.9699 0.9281 0.9485 0.9970 278
YAW 0.8113 0.9724 0.8846 0.9571 942
Overall 0.9512 2201
Accuracy 0.8274

### v3

@@ Abstract
Foundation Models, which leverage large neural networks pre-trained on unlabelled data before fine-tuning for specific tasks, are increasingly being applied to specialised domains. Recent examples include ClimaX for climate and Clay for satellite Earth observation, but a Foundation Model for Space Object Behavioural Analysis has not yet been developed. As orbital populations grow, automated methods for characterising space object behaviour are crucial for space safety. Here, we present a self-supervised framework for space object behavioural analysis, representing a first step towards a Foundation Model for SOBA. The backbone of this framework is a Perceiver-Variational Autoencoder (VAE) architecture, pre-trained with self-supervised reconstruction and masked reconstruction on 227,000 light curves from the MMT-9 observatory. The VAE enables anomaly detection, space object motion prediction, and generation of synthetic light curves. We fine-tuned the model for anomaly detection & motion prediction using two independent light curve simulators (CASSANDRA and GRIAL respectively), using CAD models of boxwing, Sentinel-3, SMOS, and Starlink platforms. Our pre-trained model achieved a reconstruction mean squared error of 0.009, identifying potentially anomalous light curves through reconstruction difficulty. After fine-tuning, the model scored 85% and 82% accuracy, with 0.92 and 0.95 ROC AUC scores respectively in both anomaly detection and motion mode prediction (e.g., sun-pointing, spin, tumbling etc.). Analysis of high-confidence anomaly predictions on real data revealed distinct patterns including characteristic object profiles and satellite glinting. The motion mode prediction model successfully differentiated between various movement behaviours such as sun-pointing, spin, and tumbling. Our work demonstrates how self-supervised learning can simultaneously enable anomaly detection, motion prediction, and synthetic data generation from rich representations learned in pre-training. More broadly, our work supports space safety and sustainability through automated monitoring and simulation capabilities.
Keywords Space Object Behavioural Analysis, Space Foundation Model, Self-Supervised Learning, Light curve anomaly detection, Attitude prediction, Space Situational Awareness (SSA), Generative AI
@@ 1 Introduction
Figure 1: Graphical structure of the paper. First, in Section 2.1 we describe pre-training our model with a large unlabelled dataset of real light curves. We encode these into a rich latent representation with a Perceiver-VAE architecture, updating these representations based on three self supervised learning (SSL) tasks: Reconstruction, Forecasting, and Masking. We next analyse the results of pre-training, which includes pre-flagging of anomalies based on reconstruction difficulty (Sections 2.2-2.3), and forecasting quality of the model (2.4). Following this, we descibe fine-tuning our rich representations for two downstream tasks: a) anomaly prediction (Section 2.5, and b) motion prediction (Section 2.6). Finally, we demonstrate further utility of our representations by generating de novo datasets according to a particular motion type (Section 2.7).
@@
However, there is little current exploration on the best neural network architectures and training strategies for FMs for more specialised datasets/applications. To our knowledge, this work represents the first step towards a Foundation Model for Space Object Behavioural Analysis. Whilst our model is pre-trained on a single modality (light curves), the Perceiver architecture is well suited to multimodal extension, and we discuss this direction in Section 4. We therefore use the term Foundation Model as a framing for this approach, following the precedent of domain-specific models such as ClimaX and Clay. Of principal interest in SDA is anomaly detection, i.e., detecting space objects that are behaving in an unusual way, such as motions, manoeuvres, or those that are performing atypical functions. Characterising these atypical behaviours can be done through simulation of space object observations, but this is challenging, due to imperfect physical models and computational cost. Therefore, general purpose models for SDA should be able to perform anomaly detection, integrating real observations with physical models in a computationally efficient way.
@@
Whilst VAEs offer an effective anomaly detection mechanism, they still need integration into a broader neural network framework to effectively process complex, high dimensional space object data. Transformer architectures are an attractive choice here, due to their ability to capture long range dependencies in sequential data. When choosing a Transformer architecture, there are numerous variants with respective strengths and weaknesses. For example, conventional transformer-based neural networks perform the computationally costly full self attention mechanism between all the inputs. This means that for very long sequences, standard transformers incur quadratic computational cost with respect to sequence length, which can become prohibitive at scale. To address this limitation, DeepMind recently developed the Perceiver/Perceiver-IO, a transformer-based neural network architecture. [14, 13]. Whilst a light-curve specific model may benefit from a state-of-the-art architecture for time-series analysis such as TimesNet [30] for anomaly detection or TimeXer [29] for forecasting, for a Space Domain Awareness Foundation Model we prioritise extensibility to other observational modalities (e.g., hyperspectral curves, ISAR images, attitude measurements, orbital parameters). To achieve this multimodal compatibility, we adopt the Perceiver architecture, which is straightforwardly extendable to multiple modalities through its cross-attention latent bottleneck. This choice also reduces the computational complexity of the vanilla transformer $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ with respect to sequence length, an advantage that is primarily forward-looking in this work but could support edge deployment (e.g., on-board a satellite).
Given the potential of VAEs as anomaly detectors, and the computational efficiency of Perceiver, a VAE-based Perceiver model is a promising approach for a computationally efficient, SDA FM for anomaly detection. In this work, we use visible light curves as a test-bed for our self-supervised framework, providing proof of principle for a SDA FM. Light curves are readily available, with large amounts of real observational data, and the maturity of light curve simulators. Here, we use the largest publicly released dataset of light curves, from the Mini-Mega TORTORA (MMT-9) observatory in Russia [4].
@@ 2 Results & Discussion
Using a held-out test portion of the MMT dataset, we examined the highest and lowest reconstruction error samples (Tables 1-2). Among the highest reconstruction error test samples (Table 1), we found a mixture of satellites and rocket bodies, some appearing frequently in the pre-training data. For example, light curves from EGS (AJISAI) occur 1064 times in the MMT dataset, yet four light curves appear in the ten highest reconstruction error. Similarly, light curves from GLOBSTAR M092 and CZ-3 R/B (occurring 285 and 269 times respectively) also showed high reconstruction errors. When an object such as AJISAI appears frequently in the pre-training data, the model learns a stable distribution of typical light curve morphologies for that object. A curve that falls outside this learned distribution is therefore more likely to represent atypical behaviour for that specific object, even if the object itself is well represented. For less frequently observed objects, the model has less prior information and cannot make the same distinction. This motivates the use of large, diverse pre-training datasets to maximise the range of objects for which such distinctions can be made. For the lowest reconstruction error samples (Table 2), we again find a mixture of satellites and rocket bodies, with one example of debris, typically with fewer appearances in the pre-training data (ranging from 9–153).
Table 1: The ten highest reconstruction errors from the pre-flagged light curve samples contained in the test set. We evaluated the lowest-loss light curve model from Section 2.2 against the entire test dataset unseen during training/validation, recording the reconstruction error for each sample. Examination of the ten highest reconstruction error light curves shows a variety of space objects (SOs). Seven light curves are from satellites (all of which are in orbit), and three belong to rocket bodies. US: United States of America, IND: India, PRC: China.
@@ 2.5 Fine-tuning for anomaly detection
Obtaining real observational data of satellites with known anomalies is extremely challenging, as the full set of possible anomalies is not well understood or catalogued, and organisations are opaque about anomalous space object behaviour for national security/commercial reasons. To address this data gap, we produced a synthetic dataset for fine-tuning using the framework CASSANDRA (Computational Agent For Space Situational Awareness And Debris Remediation Automation). Within this framework is an orbital simulator first introduced by Vasile et al. (2023), generating the light curves using a CAD model of an orbiting Space Object with a specific geometry. This dataset contains 800 simulations of four different satellite platforms: a boxwing satellite (e.g., Jason-3), Sentinel-3, SMOS, and Starlink. These space objects (SOs) orbit an observing ground station, with both light and spectral curves recorded as time courses. Within these 800 simulations, a subset contain anomalous events of varying magnitudes (specifically debris collisions that result in detectable changes in signal). These changes in signal are shown in Figure 6 F, H, where otherwise linear light curves abruptly change to a different profile, oscillating in magnitude between very high and low minima. We first evaluated the anomaly detection capability of the largest pre-trained model without any labelled data, we passed a balanced set of 32 anomalous and 32 normal CASSANDRA light curves through the frozen encoder and used reconstruction error as an anomaly score, flagging the top 10% highest errors as anomalous. This unsupervised approach achieved 61% accuracy and a ROC AUC of 0.71, demonstrating that the pre-trained representations capture anomaly-relevant structure without any supervision. Fine-tuning with labelled data substantially improves on this, as described below. To fine-tune our pre-trained Perceiver-VAE for anomaly detection, we freeze the encoder weights and train a lightweight MLP classification head on the learned latent representations (see Section 3.4 for details).
We compare against four supervised baselines trained from scratch on raw light curves, as well as randomly initialised Perceiver encoders (r.e.) at increasing scale (Table 3). Among the baselines, the CNN achieves the strongest ROC AUC (0.849), with the LSTM achieving the best accuracy (0.856). Randomly initialised Perceivers are broadly competitive with these baselines, suggesting the architecture itself has useful inductive biases for this task. Pre-training consistently improves on random initialisation across all scales, with pre-trained models achieving ROC AUC of 0.911–0.916. Our full model achieves 0.850 accuracy, 0.916 ROC AUC, and 0.793 macro F1, outperforming all baselines across metrics.
@@
To assess label efficiency, we evaluated the pre-trained Perceiver against the CNN baseline in a $k$-shot regime, training each model on $k\in\{2,5,10,25\}$ labelled examples per class across 20 independent runs (Table 4). The pre-trained Perceiver outperforms the CNN at all values of $k$, including as few as two labelled examples per class (ROC AUC $0.777\pm 0.082$ vs. $0.731\pm 0.049$). The advantage of pre-training grows with $k$, reaching a gap of $0.082$ AUC points at $k=25$. This demonstrates that the representations learned during pre-training are particularly valuable in the label-scarce regime typical of anomaly detection in operational SDA settings.
@@
Table 3: Anomaly detection performance on the held-out test set. Supervised baselines are trained from scratch on raw light curves. Perceiver models are shown at increasing scale; random encoder (r.e.) variants confirm that pretrained representations drive performance.
Model Size Accuracy ROC AUC Macro F1
RNN 0.831 0.635 0.587
LSTM 0.856 0.741 0.783
CNN 0.819 0.849 0.743
Transformer 0.800 0.774 0.672
Perceiver (r.e.) s 0.756 0.747 0.559
Perceiver (r.e.) m 0.756 0.734 0.559
Perceiver (pre.) s 0.825 0.911 0.763
Perceiver (pre.) m 0.856 0.912 0.796
Ours (pre.) l 0.850 0.916 0.793
Table 4: Label efficiency comparison for anomaly detection. ROC AUC (mean $\pm$ std. dev. across 20 independent runs) at increasing numbers of labelled training examples ($k$ per class). The pre-trained Perceiver (large) outperforms the CNN baseline at all values of $k$.
Model ROC AUC by $k$ shots per class
$k=2$ $k=5$ $k=10$ $k=25$
CNN (baseline) 0.731 $\pm$ 0.049 0.754 $\pm$ 0.034 0.754 $\pm$ 0.009 0.780 $\pm$ 0.008
Perceiver (pre.) 0.777 $\pm$ 0.082 0.800 $\pm$ 0.109 0.851 $\pm$ 0.040 0.862 $\pm$ 0.030
Having fine-tuned our model to high classification accuracy, we next analysed the held-out test dataset, classifying test samples as anomalous/non-anomalous and visualising the top twelve most confidently predicted anomalies (Figure 7). We found a variety of SOs in this set, including several satellites (Figure 7 A, B, E, H, J, L). All of these satellites except the light curve belonging to FLOCK 4P 6 (Figure 7 E) exhibited ’J’ shaped curves, which simulations suggest may represent glinting from a highly reflective surface aligning with the sun [18]. We also detected light curves belonging to debris (Figure 7 C, D, G), all displaying prominent ’U’ shaped profiles i.e., where the minima in magnitude is approximately in the middle of the observation. Finally, we identified three other distinct light curve motifs (Figure 7 red dotted outlines): abrupt changes in magnitude which presented as peaks (Figure 7 E), step changes (Figure 7 F), or highly periodic variation (Figure 7 K). We note that whilst these predictions are made on real MMT-9 observations, no ground truth anomaly labels exist for this dataset and therefore real anomalies cannot be confirmed this way. The identified morphologies are therefore interpreted qualitatively, guided by known physical signatures from the literature and simulation. Systematic validation against catalogued anomaly events or operator records remains an important direction for future work.
@@ 2.6 Fine-tuning for motion prediction
We also fine-tuned on a separate dataset with ground-truth motions, simulated by the GRIAL simulator from GMV [8]. Example light curves from this dataset are visualised in Figure 8, here we briefly summarise the dataset.
Table 5: Comparison of motion classification performance on the held-out test set. Supervised baselines are trained from scratch on raw light curves. Perceiver models are shown at increasing scale; random encoder (r.e.) variants use randomly initialised weights to confirm that pretrained representations drive performance. Our full model (end-to-end finetuning) achieves best performance.
Model Size Accuracy ROC AUC Macro F1
RNN 0.354 0.686 0.273
Transformer 0.521 0.753 0.355
LSTM 0.693 0.920 0.652
CNN 0.711 0.911 0.644
Perceiver (r.e.) s 0.471 0.625 0.199
Perceiver (r.e.) m 0.432 0.500 0.101
Perceiver (pre.) s 0.635 0.860 0.541
Perceiver (pre.) m 0.638 0.871 0.556
Perceiver (pre.) l 0.733 0.907 0.628
Ours (pre., e2e) l 0.827 0.951 0.773
@@
The GMV dataset is a simulated dataset of 22,009 light curves of Sentinel-3A simulated under 10 different motion laws as seen from 30 different ground stations. These motions can be grouped into three distinct behaviours: First, Sun oriented: SafeX, SafeZ (where the Space Object’s X/Z axis points to the Sun respectively, and Y initially points to the celestial North pole, and the X/Z rotates respectively), and Sun (X-axis points to Sun, Y-axis to Celestial North pole, differentiated from Safe by a varied X-axis phase angle. Next, Earth oriented motions: YAWXC, YAWZC, YAWXS, YAWZS (where X/Z indicates the Space Object’s nadir pointing axis, and C/S indicates whether the object’s motion is compensating (C) for optical distortion, or (S) maximising solar array lighting). Finally, the dataset also contains labelled general motions less related to satellite function: Tumbling (uncontrolled/complex rotation of the SO), Spin (rotation around a single axis), Inertial pointing (where the SO has a fixed orientation relative to the J2000 reference frame). To fine-tune our pre-trained Perceiver-VAE, we first group similar motions together, as domain expert manual inspection between e.g., different yaw axial motions is a substantially more challenging task than distinguishing between different motions altogether (e.g., tumbling/sun pointing). After fine-tuning for 100 epochs, our motion classification tuned Perceiver-VAE could classify held out light curves with 82.74% accuracy, with an average ROC-AUC of 0.95 (Table 6). There was inter-class variation in the fine-tuned classifier’s performance. For example, both tumbling and spin show high performance, with precision of approximately 0.97 and 0.90 respectively. Similarly, these classes also showed high recall, both with approximately 0.93. In contrast, inertial has high precision (0.86) but low recall (0.33). This is potentially due to morphological similarity between inertial and yaw light curves (the largest class), both of which exhibit smooth, monotonically increasing profiles without substantial local variation in magnitude (Figure 8, compare A and E). Table 5 compares our model against supervised baselines and randomly initialised Perceiver-VAEs of varying sizes. Among the baselines, the LSTM and CNN models achieve the strongest performance (ROC AUC 0.92 and 0.91 respectively), with both the RNN (ROC AUC 0.69) and Transformer (ROC AUC 0.75) weaker on this task. Similarly, randomly initialised perceiver encoders perform poorly (ROC AUC 0.63 and 0.50 for the small and medium variant respectively). Pre-training provides substantial improvements as feature extraction at all scales, with the pre-trained Perceiver-VAEs reaching 0.86-0.91 ROC AUC. Our full large model uses end-to-end fine-tuning of the pre-trained encoder alongside the classification head (see Section 3.4 for details), which further increases classification metrics (0.95 ROC AUC, with 83% accuracy and 0.77 F1 score).
@@
Table 6: Performance evaluation of the motion classification model on a held-out test dataset. The model achieves 82.74% overall accuracy with an average ROC AUC of 0.9512. The metrics demonstrate varied performance across different motion classes, with TUMBLING and SPIN showing the strongest classification performance, while INERTIAL exhibits high precision but low recall. For YAW, the most common class, the model shows high recall but lower precision. $n_{\text{train}}$ denotes the number of training samples per class.
Motion Class Precision Recall F1-Score ROC AUC n ntrain
INERTIAL 0.860 0.328 0.475 0.869 262 1,259
SAFE 0.771 0.749 0.760 0.948 342 1,940
SPIN 0.900 0.929 0.915 0.991 225 1,280
SUN 0.676 0.632 0.653 0.945 152 940
TUMBLING 0.970 0.928 0.949 0.997 278 1,333
YAW 0.811 0.972 0.885 0.957 942 5,131
Overall 0.951 2201 11,883
Accuracy 0.827